# UWPD: A General Paradigm for Invisible Watermark Detection Agnostic to Embedding Algorithms

### 06 Mar 2026 v1

@@ Abstract
Invisible watermarks, as an essential technology for image copyright protection, have been widely deployed with the rapid development of social media and AIGC. However, existing invisible watermark detection heavily relies on prior knowledge of specific algorithms, leading to limited detection capabilities for "unknown watermarks" in open environments. To this end, we propose a novel task named Universal Watermark Presence Detection (UWPD), which aims to identify whether an image carries a copyright mark without requiring decoding information. We construct the UniFreq-100K dataset, comprising large-scale samples across various invisible watermark embedding algorithms. Furthermore, we propose the Frequency Shield Network (FSNet). This model deploys an Adaptive Spectral Perception Module (ASPM) in the shallow layers, utilizing learnable frequency gating to dynamically amplify high-frequency watermark signals while suppressing low-frequency semantics. In the deep layers, the network introduces Dynamic Multi-Spectral Attention (DMSA) combined with tri-stream extremum pooling to deeply mine watermark energy anomalies, forcing the model to precisely focus on sensitive frequency bands. Extensive experiments demonstrate that FSNet exhibits superior zero-shot detection capabilities on the UWPD task, outperforming existing baseline models. Code and datasets will be released upon acceptance.
@@ 1 Introduction
The explosive development of Generative AI (AIGC) [12, 23, 26] and short-video social media has drastically lowered the threshold for image creation and dissemination, but it has simultaneously triggered an unprecedented crisis of intellectual property (IP) and copyright abuse. In this context, invisible watermarking technology has become an essential means of copyright protection due to its advantages of minimal image degradation and high resistance to tampering. With the escalation of application demands, invisible watermarking technology has undergone rapid iteration, evolving from early Least Significant Bit (LSB) steganography [31] and Discrete Cosine Transform (DCT) frequency embedding [6, 35] to recent deep learning-based [41, 29] and generative model watermarks [8, 34, 10]. However, this technological diversity has also spawned a severe problem:
The extraction of existing invisible watermarks relies on decoding algorithms that strictly correspond to their embedding methods, making "universal invisible watermark decoding" almost impossible in practical applications.
@@
Nevertheless, although universal invisible watermark decoding is difficult to achieve, certain commonalities exist among invisible watermarking technologies; namely, to ensure visual imperceptibility, invisible watermarks typically embed features implicitly into the high-frequency bands of images [38, 1, 5]. Based on this commonality, we attempt to construct a universal invisible watermark presence detection(UWPD) mechanism, serving as a Zero-Trust preliminary screening barrier to avoid infringement by identifying the existence of invisible watermarks in images. Figure 1 illustrates the background aforementioned.
Figure 1: Background and motivation of the UWPD task showing the challenges of universal watermark detection.
To address such issues, we propose a novel computer vision task: Universal Watermark Presence Detection (UWPD). Since it is impossible to ensure that the dataset or model covers all invisible watermarking technologies, the UWPD task places greater emphasis on the detection capability for unknown types of watermarks, that is, the zero-shot performance.
The training of universal models heavily relies on data diversity. Therefore, we construct a large-scale comprehensive dataset named UniFreq-100K for the UWPD task. This dataset not only contains watermarked images covering multiple image creation domains and various invisible watermark embedding algorithms, but also provides an equal scale of clean images, offering a benchmark tailored to real-world scenarios for the UWPD task.
@@
Extensive experiments and ablation analyses demonstrate that FSNet significantly outperforms existing detection baselines on the UWPD task, exhibiting superior zero-shot detection capabilities and strong robustness against unseen watermarks.In summary, the main contributions of this paper are as follows:
  • Defining a new task and benchmark: We propose the UWPD task for the first time and construct UniFreq-100K, a large-scale dataset covering multiple watermark generation paradigms, providing a novel research paradigm for digital image copyright protection.
@@
  • Validating the model’s effectiveness on the task: We evaluate the performance of mainstream vision models and FSNet on the UWPD task. Among them, FSNet demonstrates performance surpassing baseline models, establishing a new baseline for the UWPD task.
@@ 2.1 Invisible Watermarks
With the advancement of deep learning, end-to-end neural watermarking frameworks utilizing encoder-decoder architectures (e.g., HiDDeN [41]) have become mainstream, significantly improving embedding capacity and efficiency [30, 39, 40]. Recently, the explosion of Generative AI (AIGC) has catalyzed exclusive, highly heterogeneous watermarking technologies for generative models [8, 34, 10], making the construction of a unified decoder practically impossible.
@@ 2.2 Vision Foundation Models
While Vision Foundation Models (e.g., ResNet [11], ViT [7]) have significantly advanced visual forensics and zero-shot fake image detection [4, 19, 33], they suffer from an inherent "low-frequency bias" [37]. During feature extraction and spatial downsampling, these models tend to capture global semantics while discarding local high-frequency details as noise. This bias severely hinders the UWPD task: since invisible watermarks are typically concealed within these microscopic high-frequency structures, traditional models struggle to capture such extremely weak features.
@@
2.3 Existing Universal Invisible Watermark Detection
Currently, Universal Watermark Detection remains a highly challenging frontier field. Existing invisible watermark extraction heavily relies on decoders strictly corresponding to the embedding algorithms. Although in the fields of multimedia forensics and steganalysis, methods such as SRM [9] and SRNet [3] attempt to detect image steganography by extracting high-frequency noise residuals, these methods are primarily designed for traditional extremely low-capacity steganography.
Faced with modern, complex, and variable deep learning watermarks and generative watermarks (such as algorithms based on latent space perturbations), traditional steganalysis methods often exhibit severe feature drift and are difficult to generalize across algorithms. Due to the lack of large-scale benchmark datasets covering diverse modern watermarks, existing detection methods possess almost no zero-shot detection capability when facing unseen algorithms.
3 Universal Watermark Presence Detection
Given an input image $x\in\mathcal{X}\subset\mathbb{R}^{H\times W\times C}$, UWPD aims to learn a binary classification mapping model $\mathcal{F}_{\theta}$ to determine whether an invisible watermark is embedded in the image. We define the label space of this task as $\mathcal{Y}=\{0,1\}$, where $y=0$ denotes the original clean image and $y=1$ denotes the watermarked image. Therefore, the core objective of UWPD is to optimize the model parameters $\theta$ so that it can accurately perform the mapping determination:
$\hat{y}=\mathcal{F}_{\theta}(x)\in\{0,1\}$ (1)
@@ 3 Universal Watermark Presence Detection
$x_{w}=\mathcal{E}_{k}(x_{c},m),\quad\text{s.t.}\quad\mathcal{D}(x_{w},x_{c})\leq\epsilon$ (2)
@@
In real-world scenarios, the embedding algorithm may belong to an extremely large and heterogeneous algorithm family $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$. Since different algorithms will hide the watermark signal in different frequency bands or feature spaces, constructing a universal decoder to extract the watermark information is mathematically highly ill-posed. Therefore, UWPD bypasses the complex decoding process and instead focuses on a more universal degenerate task: detecting the embedding traces left by any unknown algorithm.
Traditional image classification heavily relies on the low-frequency semantic content of images, whereas to satisfy the visual imperceptibility constraint, invisible watermarks typically do not alter the macroscopic visual representation of the image, but instead hide perturbation signals in high-frequency artifacts to which the human eye is extremely insensitive. This implies that for the UWPD model, there is a need to shift the model from capturing the visible macroscopic features of the image to capturing the invisible high-frequency features.
@@ 4.1 Dataset Description
Based on this commonality, we employ representative sampling, selecting representative algorithms from existing mainstream watermarking paradigms, ranging from early spatial and frequency domains to modern deep learning and generative models, to construct the UWPD benchmark dataset—UniFreq-100K.
@@
Meanwhile, the dataset also contains an equal amount of unwatermarked samples covering multiple sources. The ratio of positive to negative samples for each source category is kept close, ensuring sample balance. Ultimately, UniFreq-100K contains 94,000 watermarked images (positive samples) and 96,000 unwatermarked images (negative samples).
Table 1: Detailed distribution of image types and watermark algorithms in the dataset. All numerical values are presented in thousands (k). Citations for the evaluated watermarking algorithms: LSB [31], Patchwork [2], DCT [6], DWT [35], HiDDeN [41], StegaStamp [30], Stable Sig. [8], Tree-Ring [34], and SynthID [10].
Category Unwat. Watermarked Algorithms Sub. Total
LSB Patch. DCT DWT HiDDeN Stega. Stable Tree Synth.
Real Photos 30 4 4 4 4 4 4 0 0 0 24 54
2D/Digital Art 12 2 2 2 2 2 2 0 0 0 12 24
AIGC Images 30 2 2 2 2 2 2 10 10 2 34 64
E-Posters 12 2 2 2 2 2 2 0 0 0 12 24
Scanned Drawings 12 2 2 2 2 2 2 0 0 0 12 24
Total 96 12 12 12 12 12 12 10 10 2 94 190
@@ 4.2 Evaluation Recommendations
The core challenge and practical significance of the UWPD task lie in the highly robust recognition of unseen watermarks. If a conventional random split is merely performed on a mixed dataset, the model is highly prone to degenerating into overfitting specific known algorithm features, thereby deviating from the original intention of universal presence detection.
To this end, we abandon the traditional train/test splitting method and propose using leave-one-algorithm-out cross-validation as the standard evaluation protocol for the UWPD task.
@@
Furthermore, to avoid threshold shift caused by class imbalance in binary classification tasks, during each round of training and testing, we sample unwatermarked images ($y=0$) from the homologous clean image library strictly matching the number of watermarked images at a 1:1 ratio, to form a complete binary classification evaluation data stream.
@@ 5 FSNet
Common deep learning backbone networks (such as ResNet, ViT, etc.) are often designed to extract low-frequency visible information, and can easily treat high-frequency watermark signals as "noise" and discard them during downsampling. To address the models’ deficiency in capturing high-frequency signals, we propose the Frequency Shield Network (FSNet), a spatial-frequency joint perception architecture specifically designed for UWPD, aiming to explicitly capture and retain spectral anomalies.
As illustrated in Figure 2, FSNet follows a Dual-Stage spectral Pperception paradigm. We do not treat frequency-domain analysis as an isolated preprocessing step, but rather deeply integrate it into the feature extraction hierarchy. This architecture consists of three core components:
@@
    Adaptive Spectral Perception Module (ASPM): Deployed at the shallowest layer (Stem) of the network, this module utilizes learnable frequency masks as "soft gating," allowing high-frequency signals related to watermarks to pass while suppressing the dominant low-frequency semantic content.
@@
$\bm{F}_{stem}=\Phi_{ASPM}(\bm{X})$ (3)
$\bm{F}_{deep}=\Phi_{Backbone}(\bm{F}_{stem})$ (4)
$\bm{Z}=\Phi_{DMSA}(\bm{F}_{deep})$ (5)
@@ 5.1 Adaptive Spectral Perception Module (ASPM)
Existing frequency-domain object detection methods typically employ fixed high-pass filters to separate edges and detail information. However, the attacks and defenses of invisible watermarks are extremely diverse. To resist corruption such as JPEG compression, many modern watermarking algorithms do not exist solely at fixed frequencies; thus, a fixed cutoff frequency would lead to a "one-size-fits-all" drawback, easily mistakenly deleting crucial watermark signals.
@@
$\hat{\bm{X}}_{freq}=\bm{X}_{freq}\odot\mathcal{I}(\bm{M})$ (7)
@@
$\bm{Y}_{ASPM}=\text{Conv}_{3\times 3}\Big(\bm{X}+\bm{R}_{spatial}+\text{MaxPool}(\bm{R}_{spatial})\Big)$ (8)
@@ 5.2 Dynamic Multi-Spectral Attention (DMSA)
After downsampling and semantic extraction by the backbone network, deep features often face information aliasing, requiring recalibration along the channel dimension (channel-wise) to distinguish watermark features from normal image textures. Inspired by multi-band analysis [18], we design the DMSA module. Unlike traditional channel attention mechanisms (such as SE-Block) that heavily rely on global average pooling (GAP), DMSA utilizes DCT basis functions to losslessly compress spatial information into spectral descriptors of specific frequencies.
@@
$B_{h,w}^{u,v}=\cos\left(\frac{\pi h}{H_{b}}(u+0.5)\right)\cos\left(\frac{\pi w}{W_{b}}(v+0.5)\right)$ (9)
To address the issue of deep-layer resolution changes in large models (such as ViT), we do not forcibly pool the input features to a fixed $7\times 7$ size as conventional approaches do. Instead, we dynamically interpolate and align the generated basis function weights, thereby maximally preserving the high-frequency fragments in deep feature maps.
Tri-Stream Extremum Pooling Mechanism: Modern invisible watermarks often employ local embedding to reduce visual degradation. The presence of watermarks manifests not only as "peaks" of frequency energy but often also as abnormal "valleys" of local frequency energy (especially in some pixel-subtraction-based embedding algorithms). Traditional average pooling or max pooling can easily miss these covert signals. Therefore, we introduce a tri-stream extremum pooling branch in DMSA.
@@
$\bm{v}_{avg}=\text{AvgPool}(\bm{S}),\quad\bm{v}_{max}=\text{MaxPool}(\bm{S}),\quad\bm{v}_{min}=-\text{MaxPool}(-\bm{S})$ (10)
@@
$\bm{v}_{total}=\sigma\left(\text{MLP}\Big(\frac{\bm{v}_{avg}+\bm{v}_{max}+\bm{v}_{min}}{K}\Big)\right)$ (11)
This weight $\bm{v}_{total}$ is used for channel re-weighting of deep features. This mechanism endows the model with "frequency band screening" capabilities, enabling it to dynamically suppress channels dominated by low-frequency semantics and assign higher activation values to channels harboring watermark spectra.
Tri-Stream Extremum Pooling: Modern invisible watermarks often employ local embedding. The presence of watermarks manifests not only as "peaks" of frequency energy but often also as abnormal "valleys" of local frequency energy. Traditional average pooling or max pooling can easily miss these covert signals. Therefore, we introduce a tri-stream extremum pooling branch (average, max, and min extremum pooling) in DMSA.
The final channel attention weight is aggregated through a multi-layer perceptron and used for channel re-weighting of deep features. This mechanism endows the model with "frequency band screening" capabilities, dynamically suppressing channels dominated by low-frequency semantics and assigning higher activation values to channels hiding watermark spectra.

### 13 Mar 2026 v2

# AWPD: Frequency Shield Network for Agnostic Watermark Presence Detection

@@ Abstract
Invisible watermarks, as an essential technology for image copyright protection, have been widely deployed with the rapid development of social media and AIGC. However, existing invisible watermark detection heavily relies on prior knowledge of specific algorithms, leading to limited detection capabilities for “unknown watermarks” in open environments. To this end, we propose a novel task named Agnostic Watermark Presence Detection (AWPD), which aims to identify whether an image carries a copyright mark without requiring decoding information. We construct the UniFreq-100K dataset, comprising large-scale samples across various invisible watermark embedding algorithms. Furthermore, we propose the Frequency Shield Network (FSNet). This model deploys an Adaptive Spectral Perception Module (ASPM) in the shallow layers, utilizing learnable frequency gating to dynamically amplify high-frequency watermark signals while suppressing low-frequency semantics. In the deep layers, the network introduces Dynamic Multi-Spectral Attention (DMSA) combined with tri-stream extremum pooling to deeply mine watermark energy anomalies, forcing the model to precisely focus on sensitive frequency bands. Extensive experiments demonstrate that FSNet exhibits superior zero-shot detection capabilities on the AWPD task, outperforming existing baseline models. Code and datasets will be released upon acceptance.
@@ 1 Introduction
The explosive development of Generative AI (AIGC) [12, 23, 26] and short-video social media platforms has democratized image creation and dissemination, yet this accessibility has simultaneously precipitated an unprecedented crisis in intellectual property (IP) protection and copyright compliance. In this context, invisible watermarking technology has become an essential means of copyright protection due to its advantages of minimal image degradation and high resistance to tampering. With the escalation of application demands, invisible watermarking technology has undergone rapid iteration, evolving from early Least Significant Bit (LSB) steganography [31] and Discrete Cosine Transform (DCT) frequency embedding [6, 35] to recent deep learning-based [42, 30] and generative model watermarks [8, 34, 10]. However, this technological diversity has also spawned a severe problem:
The extraction of existing invisible watermarks relies on decoding algorithms that strictly correspond to their embedding methods, making “agnostic invisible watermark decoding” almost impossible in practical applications.
@@
Nevertheless, although agnostic invisible watermark decoding is difficult to achieve, certain commonalities exist among invisible watermarking technologies; namely, to ensure visual imperceptibility, invisible watermarks typically embed features implicitly into the high-frequency bands of images [39, 1, 5]. Based on this commonality, we attempt to construct a agnostic invisible watermark presence detection(AWPD) mechanism, serving as a Zero-Trust preliminary screening barrier to avoid infringement by identifying the existence of invisible watermarks in images.
To address such issues, we propose a novel computer vision task: Agnostic Watermark Presence Detection (AWPD). Since it is impossible to ensure that the dataset or model covers all invisible watermarking technologies, the AWPD task places greater emphasis on the detection capability for unknown types of watermarks, that is, the zero-shot performance.
The training of agnostic models heavily relies on data diversity. Therefore, we construct a large-scale comprehensive dataset named UniFreq-100K for the AWPD task. This dataset not only contains watermarked images covering multiple image creation domains and various invisible watermark embedding algorithms, but also provides an equal scale of clean images, offering a benchmark tailored to real-world scenarios for the AWPD task.
@@
Extensive experiments and ablation analyses demonstrate that FSNet significantly outperforms existing detection baselines on the AWPD task, exhibiting superior zero-shot detection capabilities and strong robustness against unseen watermarks.In summary, the main contributions of this paper are as follows:
  • Defining a new task and benchmark: We propose the AWPD task for the first time and construct UniFreq-100K, a large-scale dataset covering multiple watermark generation paradigms, providing a novel research paradigm for digital image copyright protection.
@@
  • Validating the model’s effectiveness on the task: We evaluate the performance of mainstream vision models and FSNet on the AWPD task. Among them, FSNet demonstrates performance surpassing baseline models, establishing a new baseline for the AWPD task.
@@ 2.1 Invisible Watermarks
With the advancement of deep learning, end-to-end neural watermarking frameworks utilizing encoder-decoder architectures (e.g., HiDDeN [42]) have become mainstream, significantly improving embedding capacity and efficiency [29, 40, 41]. Recently, the explosion of Generative AI (AIGC) has catalyzed exclusive, highly heterogeneous watermarking technologies for generative models [8, 34, 10], making the construction of a unified decoder practically impossible.Furthermore, sophisticated defense mechanisms increasingly rely on subtle imaging perturbations rather than conventional watermarking. For instance, recent works have demonstrated that injecting imperceptible adversarial noise to perturb attention mechanisms can efficiently fool customized diffusion models [37], which further exacerbates the complexity of detecting artificial traces in AIGC imagery.
@@ 2.2 Vision Foundation Models
While Vision Foundation Models (e.g., ResNet [11], ViT [7]) have significantly advanced visual forensics and zero-shot fake image detection [4, 19, 33], they suffer from an inherent “low-frequency bia” [38]. During feature extraction and spatial downsampling, these models tend to capture global semantics while discarding local high-frequency details as noise. This bias severely hinders the AWPD task: since invisible watermarks are typically concealed within these microscopic high-frequency structures, traditional models struggle to capture such extremely weak features.
@@
2.3 Existing Agnostic Invisible Watermark Detection
Currently, agnostic Watermark Detection remains a highly challenging frontier. Existing invisible watermark extraction heavily relies on decoders strictly corresponding to the embedding algorithms. Although in the fields of multimedia forensics and steganalysis, methods such as SRM [9] and SRNet [3] attempt to detect image steganography by extracting high-frequency noise residuals, these methods are primarily designed for traditional extremely low-capacity steganography.
Faced with modern, complex, and diverse deep learning watermarks and generative watermarks (such as algorithms based on latent space perturbations), traditional steganalysis methods often exhibit severe feature drift and are difficult to generalize across algorithms. Due to the lack of large-scale benchmark datasets covering diverse modern watermarks, existing detection methods possess almost no zero-shot detection capability when facing unseen algorithms.
Figure 1: Overall distribution of the UniFreq-100K dataset. (a) Distribution across five image categories totaling 190K images. (b) Distribution across watermarking algorithms, including 96K unwatermarked and 94K watermarked images. Evaluated watermarking algorithms: LSB [31], Patchwork [2], DCT [6], DWT [35], HiDDeN [42], StegaStamp [29], Stable Sig. [8], Tree-Ring [34], and SynthID [10]. Detailed cross-distribution data (category $\times$ algorithm) can be found in the Appendix.
3 Agnostic Watermark Presence Detection
Given an input image $x\in\mathcal{X}\subset\mathbb{R}^{H\times W\times C}$, AWPD aims to learn a binary classification mapping model $\mathcal{F}_{\theta}$ to determine whether an invisible watermark is embedded in the image. We define the label space of this task as $\mathcal{Y}=\{0,1\}$, where $y=0$ denotes the original clean image and $y=1$ denotes the watermarked image. Therefore, the core objective of AWPD is to optimize the model parameters $\theta$ so that it can accurately perform the mapping determination:
$\hat{y}=\mathcal{F}_{\theta}(x)\in\{0,1\}.$ (1)
@@ 3 Agnostic Watermark Presence Detection
$x_{w}=\mathcal{E}_{k}(x_{c},m),\quad\text{s.t.}\quad\mathcal{D}(x_{w},x_{c})\leq\epsilon,$ (2)
@@
In real-world scenarios, the embedding algorithm may belong to an extremely large and heterogeneous algorithm family $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$. Since different algorithms will hide the watermark signal in different frequency bands or feature spaces, constructing a agnostic decoder to extract the watermark information is mathematically highly ill-posed. Therefore, AWPD bypasses the complex decoding process and instead focuses on a more agnostic degenerate task: detecting the embedding traces left by any unknown algorithm.
While traditional image classification relies heavily on low-frequency semantics, invisible watermarks preserve macroscopic visual representations to satisfy imperceptibility constraints. Consequently, they conceal perturbation signals within high-frequency artifacts, rendering them imperceptible to the human visual system. This implies that for the AWPD model, there is a need to shift the model from capturing the visible macroscopic features of the image to capturing the invisible high-frequency features.
@@ 4.1 Dataset Description
Based on this commonality, we employ representative sampling by selecting algorithms from existing mainstream watermarking paradigms, ranging from early spatial and frequency domains to modern deep learning and generative models, to construct the AWPD benchmark dataset—UniFreq-100K.
@@
Meanwhile, the dataset also contains an equal amount of unwatermarked samples covering multiple sources. The ratio of positive to negative samples for each source category is kept close, ensuring sample balance. Ultimately, UniFreq-100K contains 94,000 watermarked images (positive samples) and 96,000 unwatermarked images (negative samples), as illustrated in the distribution charts Figure 1.
@@ 4.2 Evaluation Recommendations
The core challenge and practical significance of the AWPD task lie in the highly robust recognition of unseen watermarks. If a conventional random split is applied to a mixed dataset, the model is highly prone to overfitting specific known algorithm features, thereby deviating from the original intention of agnostic presence detection.
To this end, we abandon the traditional train/test splitting method and propose using leave-one-algorithm-out cross-validation as the standard evaluation protocol for the AWPD task.
@@
Furthermore, to avoid threshold shift caused by class imbalance, during each round of training and testing, we sample unwatermarked images ($y=0$) from the homologous clean image library to strictly match the number of watermarked images at a 1:1 ratio, to form a complete binary classification evaluation data stream.
@@ 5 FSNet
Standard deep learning backbones (e.g., ResNet, ViT) are typically optimized for low-frequency visible information, often treating high-frequency watermark signals as “noise” and prematurely discarding them during downsampling. To address the models’ deficiency in capturing high-frequency signals, we propose the Frequency Shield Network (FSNet), a spatial-frequency joint perception architecture specifically designed for AWPD, aiming to explicitly capture and retain spectral anomalies.
As illustrated in Figure 2, FSNet follows a Dual-Stage spectral perception paradigm. We do not treat frequency-domain analysis as an isolated preprocessing step, but rather deeply integrate it into the feature extraction hierarchy. This architecture consists of three core components:
@@
    Adaptive Spectral Perception Module (ASPM): Deployed at the shallowest layer (Stem) of the network, this module utilizes learnable frequency masks as ”soft gating,” allowing high-frequency signals related to watermarks to pass while suppressing the dominant low-frequency semantic content.
@@
$\bm{F}_{stem}=\Phi_{ASPM}(\bm{X}),$ (3)
$\bm{F}_{deep}=\Phi_{Backbone}(\bm{F}_{stem}),$ (4)
$\bm{Z}=\Phi_{DMSA}(\bm{F}_{deep}),$ (5)
@@ 5.1 Adaptive Spectral Perception Module (ASPM)
Existing frequency-domain object detection methods typically employ fixed high-pass filters to separate edges and detail information. However, the attacks and defenses of invisible watermarks are extremely diverse. To resist corruptions like JPEG compression, modern watermarking algorithms distribute signals across varying frequencies. Consequently, applying a fixed cutoff frequency introduces a rigid bottleneck, risking the unintended deletion of crucial watermark clues.
@@
$\hat{\bm{X}}_{freq}=\bm{X}_{freq}\odot\mathcal{I}(\bm{M}),$ (7)
@@
$\bm{Y}_{ASPM}=\text{Conv}_{3\times 3}\Big(\bm{X}+\bm{R}_{spatial}+\text{MaxPool}(\bm{R}_{spatial})\Big).$ (8)
@@ 5.2 Dynamic Multi-Spectral Attention (DMSA)
After downsampling and semantic extraction by the backbone network, deep features often suffer from information aliasing, requiring recalibration along the channel dimension (channel-wise) to distinguish watermark features from normal image textures. Inspired by multi-band analysis [18], we design the DMSA module. Unlike traditional channel attention mechanisms (such as SE-Block) that heavily rely on global average pooling (GAP), DMSA utilizes DCT basis functions to losslessly compress spatial information into spectral descriptors of specific frequencies.
@@
$B_{h,w}^{u,v}=\cos\left(\frac{\pi h}{H_{b}}(u+0.5)\right)\cos\left(\frac{\pi w}{W_{b}}(v+0.5)\right).$ (9)
To address the issue of deep-layer resolution changes in large models (such as ViT), we avoid forcibly pooling the input features to a fixed $7\times 7$ size, unlike conventional approaches. Instead, we dynamically interpolate and align the generated basis function weights, thereby maximally preserving the high-frequency fragments in deep feature maps.
Tri-Stream Extremum Pooling Mechanism: Modern invisible watermarks often employ local embedding to reduce visual degradation. The presence of watermarks manifests not only as ”peaks” of frequency energy but often also as abnormal ”valleys” of local frequency energy (especially in some pixel-subtraction-based embedding algorithms). Traditional average pooling or max pooling can easily miss these covert signals. Therefore, we introduce a tri-stream extremum pooling branch in DMSA.
@@
$\bm{v}_{avg}$ $=\text{AvgPool}(\bm{S}),$ (10)
$\bm{v}_{max}$ $=\text{MaxPool}(\bm{S}),$
$\bm{v}_{min}$ $=-\text{MaxPool}(-\bm{S}),$
@@
$\bm{v}_{total}=\sigma\left(\text{MLP}\Big(\frac{\bm{v}_{avg}+\bm{v}_{max}+\bm{v}_{min}}{K}\Big)\right).$ (11)
This weight $\bm{v}_{total}$ is used for channel re-weighting of deep features. This mechanism endows the model with ”frequency band screening” capabilities, enabling it to dynamically suppress channels dominated by low-frequency semantics and assign higher activation values to channels harboring watermark spectra.